# Position: The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### v1

@@ 3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik & Karbasi, 2025; Zhou et al., 2025; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman & Panickssery, 2025). Richer CoT also increase privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as intermediate reasoning grows (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025b). Our practice is to cap best-of-$N$ and chain-of-thought length at evaluation and to filter or redact private content in traces (see section 6 (C)), while reporting many-shot jailbreak stress tests and a clear refusal policy as part of standard evaluation (see section 6 (D)).
@@
These taxes distort reported gains in predictable ways: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability—compose rewards so correctness, grounding, and abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is priced in (section 4, section 6).
4 Perspective II — Pitfalls in Evaluation: Are We Measuring Real Reasoning Gains Accurately?
Claims about RLVR progress are highly sensitive to sampling budgets, metric design, and dataset hygiene. This matters for our thesis: budget mismatch, fragile judge pipelines, and calibration drift can overstate “reasoning gains” and complicate transfer beyond math and code. We treat contamination separately in Section 5 and focus here on budgeting, metric robustness, and calibration.
@@ 4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Wen et al., 2025a; Wu et al., 2025b). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean ± std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.
@@ 4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones—e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction. In those cases, we fall back to LLM-as-a-judge. LLM-judge pipelines offer convenience but can be brittle: small changes—seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings—produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025b). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024a). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
@@ 4.3 Calibration Drift and Overconfidence
RL optimization often sharpens the output distribution. Top-1 accuracy can rise even as calibration worsens, increasing brittleness under distribution shift (Hochlehnert et al., 2025). Combined with the volatility noted above (Sun et al., 2025a), some apparent “wins” likely reflect confident exploitation of evaluation quirks rather than robust reasoning. Evaluations should therefore track expected calibration error (ECE), output entropy, and refusal/abstention alongside accuracy, and consider early stopping or annealing when calibration degrades even if reward continues to increase. See section 6(D) for our unified checklist.
Table 2: Reported scores vs. standardized evaluation (avg@32 of pass@1, estimated by averaging 32 independent single-sample decodes) with matched decoding budgets and a shared verifier/prompt family. $\Delta$ denotes Reported - Standardized. See Appendix 4 for comprehensive standardized results of recent RLVR models.
@@
Open-RS3-1.5B AIME-24 46.70 (pass@1; Dang & Ngo 2025) 30.94 +15.76+15.76
@@
Polaris-7B-Preview AIME-24 72.60 (avg@32; NLP & collaborators 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; NLP & collaborators 2025) 93.59 4.59-4.59
@@
To make these issues concrete, we compared widely cited checkpoints to parity-controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions. The gaps in Table 2 arise from three main factors:
Sampling budgets. Multi-sample reporting (e.g., avg@64 with T=0.6T{=}0.6, top-p=0.95p{=}0.95 for STILL-3-1.5B) inflates scores relative to single-shot pass@1 under the same verifier (Min et al., 2025). In our standardized runs this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (NLP & collaborators, 2025).
Template and version drift. Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (NLP & collaborators, 2025; Liu & contributors, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
Metric and decision rules. Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025b).
@@
Under unmatched budgets, unprobed judges, and untracked calibration, distributional sharpening can masquerade as capability expansion. Clean, parity-controlled evaluation often shrinks or flips celebrated gains. Hence our checklist—budget parity and saturation curves, variance disclosure, judge robustness, calibration metrics, and a contamination audit (Section 5)—is essential to distinguish real progress from artifacts and to make the tax-aware protocol credible.
5 Perspective III — Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Liu et al., 2024b) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math.
Table 3: Partial–prompt contamination summary. The Math block (left) reports answer–match accuracy at an 80% prefix on two legacy sets vs. the fresh AIME–2025. The SimpleQA block (right) is a non–math control, where Qwen’s advantage largely attenuates. Full results, including ACC@60/40 and ROUGE–L/EM for all models and datasets, are in Tables 610 in Appendix.
@@ 5 Perspective III — Data Contamination
MATH-500 AMC-23 AIME-2025 R@80 EM@80
Llama-3-1.8B 2.8 0.0 0.0 37.11 19.86
Qwen2.5-Math-7B 58.0 52.5 0.0 29.34 12.47
Qwen2.5-32B 60.0 52.5 3.3 41.06 23.09
Qwen3-14B-Base 58.2 47.5 0.0 37.24 19.40
@@
Following Wu et al. (2025b), we reveal only the first $x\%$ of each problem ($x\!\in\!\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@x, EM@x, and ACC@x. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
@@
As summarized in Table 3, Qwen2.5–Math–7B and Qwen3–14B–Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME–2025 (0–3%). Llama–3–1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern—strong on older math sets, absent on the newest release—aligns with contamination in Qwen pretraining and clean evaluation for Llama. (See App. Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar “contamination-free” stance and methodology are advocated in code by Liu et al. (2024b) and in broader evaluations by Liang et al. (2023).
@@
On a non–math control (SimpleQA) there is no systematic Qwen advantage: for example, Qwen2.5–32B attains 41.06/23.09 (R@80/EM@80) versus Llama–3–1.8B at 37.11/19.86 (App. Table 10). The attenuation on an unseen domain supports the interpretation that elevated partial–prompt math scores on legacy sets reflect contamination rather than a general suffix–reconstruction ability.
@@
6 A Tax-Aware RLVR Protocol
Guided by the RLVR “tax” framing (§3), the evaluation pitfalls analysis (§4), and the contamination audit (§5), we specify a tax-aware protocol that preserves RLVR gains while controlling hidden costs. It is designed to curb failure modes yet remain easy to adopt in current pipelines.
(A) Separate what you care about in the reward.
We use componentized rewards that explicitly score three things: answer correctness, grounding/citation sufficiency, and calibrated refusal when evidence is weak. In practice, models train more stably when these components are scheduled, not blended from the start: learn to answer first, then add grounding, and finally layer in refusal once accuracy plateaus or overconfidence rises (Sim et al., 2025; Ren et al., 2025). If a learned judge is unavoidable, measure its fragility via small prompt or format perturbations and “one-token” hacks and report robustness deltas (Zhao et al., 2025b).
(B) Control variance and match difficulty to capacity.
Concentrate exploration at high-entropy forks early, then anneal (Cheng et al., 2025a; Agarwal et al., 2025; Cui et al., 2025; Wang et al., 2025a). Reduce waste via selective sampling: filter trivial or unsolvable prompts and down-sample near-deterministic rollouts (Xiong et al., 2025; Xu et al., 2025; Zheng et al., 2025). Use light negative reinforcement to suppress spurious high-reward modes (Zhu et al., 2025). Match difficulty to capacity by switching between imitation and GRPO via an online hardness signal (e.g., GHPO) (Liu et al., 2025d). Monitor KL to the base model, sequence entropy, and reward variance, and cool down or early-stop on variance collapse or KL spikes.
(C) Mix offline and online; stop by calibration, not reward alone.
Strong offline rollouts are an efficient bootstrap; we switch to online sampling when the offline set saturates. Crucially, we stop or cool down by calibration rather than reward alone: we track expected calibration error (ECE), output entropy, and refusal rates alongside accuracy, and we anneal or halt when ECE worsens or abstention collapses even if reward continues to rise (Hochlehnert et al., 2025; Leng et al., 2024). For deployment, we cap best-of-$N$ attempts and chain-of-thought length, and we consider “think-when-needed” policies to avoid unnecessary long reasoning traces that expand the attack and privacy surface (Fang et al., 2025).
(D) Evaluation and reporting are part of the protocol, not an afterthought.
This protocol standardizes five practices: (1) match sampling budgets for base and RLVR and publish saturation curves (accuracy vs. $k$) with area-under-curve summaries, avoiding pass@k apples-to-oranges comparisons; (2) disclose variance ($\geq$3 seeds with 95% CIs) and the decoding/TP/DP settings that can move results (Jordan et al., 2020; Patterson et al., 2024); (3) track calibration and abstention (ECE, entropy, refusal policy/rates) alongside accuracy (Hochlehnert et al., 2025; Leng et al., 2024); (4) when an LLM judge is used, report judge robustness under prompt/format perturbations and adversarial sweeps (Zhao et al., 2025b); and (5) audit contamination via fuzzy/lexical matches and partial-prompt reconstruction (ACC/EM/ROUGE at 80/60/40%), including a clean held-out set (e.g., AIME-2025) before claiming reasoning gains (Wu et al., 2025b).
(E) A compact default recipe.
In code, we instantiate the reward as
$\mathcal{R}=\alpha\,\text{Correct}+\beta\,\text{Grounding}+\gamma\,\text{Refusal}$
starting with correctness only, adding grounding midway, and introducing refusal once accuracy stalls or ECE climbs. Exploration begins hot and cools (higher temperature or entropy bonus $\rightarrow$ lower), with token-local entropy targeting at uncertain tokens and down-sampling of near-deterministic rollouts; optional negative sampling is applied when spurious high-reward modes appear (Zhu et al., 2025). Objective selection switches online between imitation and GRPO using a running hardness signal (small-$k$ pass@k or semantic entropy) (Liu et al., 2025d; Chen et al., 2025b). A calibration gate halts or anneals when ECE exceeds the baseline by a fixed margin or refusal falls below a floor—even if reward improves. The evaluation pack includes matched budgets and saturation curves, $\geq$3 seeds with CIs, calibration/entropy/refusal metrics, judge robustness deltas (if used), and a contamination audit with a clean out-of-distribution set (Hochlehnert et al., 2025; Sun et al., 2025a; Wu et al., 2025b).
(F) Where this already works in the wild.
Recent mechanisms illustrate the spirit of this protocol. ProRL uses prolonged RL with KL control and periodic resets to broaden support—effective, but calibration-sensitive (Liu et al., 2025a). He et al. (2025a) add unlikeliness rewards to counter rank bias and escape distribution sharpening. GHPO adapts the objective by difficulty to stabilize training (Liu et al., 2025d). Light-R1 shows that curricula spanning SFT/DPO/RL reduce variance and overclaiming (Wen et al., 2025a). SoberScore enforces budget parity and normalization, shrinking prominent gains (Hochlehnert et al., 2025). ThinkLess trims unnecessary chain-of-thought, reducing safety and privacy exposure with minor accuracy cost (Fang et al., 2025). And contamination audits via partial-prompt probes help distinguish memorization from reasoning (Wu et al., 2025b).
We encourage adopting these choices to re-anchor RLVR to reliability: they keep the practical wins while reducing the degrees of freedom through which budget, metric, and contamination artifacts can masquerade as genuine reasoning gains.

### v2

@@ 3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik and Karbasi, 2025; Zhou et al., 2025a; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman and Panickssery, 2025). Richer CoT also increases privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as intermediate reasoning grows (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025b).
Threat Model (Reasoning Traces). Who sees traces: exposure risk is highest for open-weight models or deployments that return full CoT; API-only hidden traces lower user-facing risk, while internal logs can aid auditing. Training vs. inference: RLVR increases trace length and attempt budgets during training and (unless capped) inference, which can amplify leakage/jailbreak surfaces. What “risk increase” means: higher probability that sensitive prompt/data fragments or unsafe behaviors appear in controllable intermediate text. When traces help: hidden/internal traces improve monitorability and post-hoc safety audits, so the goal is controlled exposure, not blanket removal.
@@
These taxes distort reported gains predictably: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability: compose rewards so correctness, grounding, and abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is accounted for in section 4 and section 6.
4 Perspective II: Pitfalls in Evaluation: Are We Measuring Real Reasoning Gains Accurately?
Claims about RLVR progress are highly sensitive to sampling budgets, metric design, and dataset hygiene. Budget mismatch, fragile judge pipelines, and calibration drift can overstate “reasoning gains” and complicate transfer beyond math and code. We treat contamination separately in section 5 and focus here on budgeting, metric robustness, and calibration.
@@ 4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Wen et al., 2025a). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean $\pm$ std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.
@@ 4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones (e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction). The evaluation often relies on LLM judges. LLM-judge pipelines offer convenience but can be brittle: small changes (seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings) produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024a). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
@@ 4.3 Calibration Drift and Overconfidence
RL optimization often sharpens the output distribution. Top-1 accuracy can rise even as calibration worsens, increasing brittleness under distribution shift (Hochlehnert et al., 2025). Combined with the volatility noted above (Sun et al., 2025a), some apparent “wins” likely reflect confident exploitation of evaluation quirks rather than robust reasoning. Evaluations should therefore track expected calibration error (ECE), output entropy, and refusal/abstention alongside accuracy, and consider early stopping or annealing when calibration degrades even if reward continues to increase. Unless otherwise stated, numeric confidence is obtained from logit-derived probabilities under the model, computed from the normalized likelihood of the extracted final answer under the same decoding setup.
Table 2: Reported scores vs. standardized evaluation (avg@32 of pass@1, estimated by averaging 32 independent single-sample decodes) with matched decoding budgets and a shared verifier/prompt family. $\Delta$ denotes Reported - Standardized. See Table 4 in the Appendix for comprehensive standardized results of recent RLVR models.
@@
Open-RS3-1.5B AIME-24 46.70 (pass@1; Dang and Ngo 2025) 30.94 +15.76+15.76
@@
Polaris-7B-Preview AIME-24 72.60 (avg@32; NLP and collaborators 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; NLP and collaborators 2025) 93.59 4.59-4.59
@@
To make these issues concrete, we compared widely cited checkpoints to budget parity controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions.
Important: some sources report pass@1 while others report avg@64 or alternative aggregation (e.g., majority vote) under different decoding settings; in those cases, $\Delta$ conflates metric and budget differences with genuine policy differences. We therefore treat Table 2 as evidence of evaluation sensitivity. Furthermore, the gaps in Table 2 are exacerbated by three main factors:
Sampling budgets.
Multi-sample reporting (e.g., avg@64 with T=0.6T{=}0.6, top-p=0.95p{=}0.95 for STILL-3-1.5B) inflates scores relative to single-shot pass@1 under the same verifier (Min et al., 2025). In our standardized runs, this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (NLP and collaborators, 2025).
Template and version drift.
Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (NLP and collaborators, 2025; Liu and contributors, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
Metric and decision rules.
Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025).
@@
Under unmatched budgets, unprobed judges, and untracked calibration, distributional sharpening can masquerade as capability expansion. Clean, parity-controlled evaluation often shrinks or flips celebrated gains. Hence, our minimum standard (budget parity and saturation curves, variance disclosure, judge robustness, calibration metrics, and a contamination audit in section 5) is necessary to distinguish real gains from artifacts.
5 Perspective III: Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Liu et al., 2024b) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math. Partial-prompt reconstruction is a high-precision but not exhaustive signal of overlap: we treat it as one useful probe to be triangulated with lexical and fuzzy matching and with fresh test sets.
Table 3: Partial-prompt contamination summary. The Math block (left) reports answer-match accuracy at an 80% prefix on two legacy sets vs. the fresh AIME-2025. The SimpleQA block (right) is a non-math control, where Qwen’s advantage largely attenuates. Full results, including ACC@60/40 and ROUGE-L/EM for all models and datasets, are in Tables 610 in the Appendix.
@@ 5 Perspective III: Data Contamination
MATH-500 AIME-2025 R@80 EM@80
Llama-3-1.8B 2.8 0.0 37.11 19.86
Qwen2.5-Math-7B 58.0 0.0 29.34 12.47
Qwen2.5-32B 60.0 3.3 41.06 23.09
Qwen3-14B-Base 58.2 0.0 37.24 19.40
@@
Following Wu and others (2025), we reveal only the first $x\%$ of each problem ($x\in\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@$x$, EM@$x$, and ACC@$x$. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
@@
As summarized in Table 3, Qwen2.5-Math-7B and Qwen3-14B-Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME-2025 (0–3%). Llama-3-1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern (strong on older math sets, absent on the newest release) is consistent with substantial overlap between legacy math benchmarks and the training corpora of some Qwen-family models, and with comparatively cleaner evaluation for Llama. (See Appendix Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar contamination-aware stance and methodology are advocated in code by Liu et al. (2024b) and in broader evaluations by Liang et al. (2023).
@@
On a non-math control (SimpleQA) there is no systematic Qwen advantage: for example, Qwen2.5-32B attains 41.06/23.09 (R@80/EM@80) versus Llama-3-1.8B at 37.11/19.86 (Appendix Table 10). The attenuation on an unseen domain supports the interpretation that elevated partial-prompt math scores on legacy sets reflect contamination rather than a general suffix-reconstruction ability. We use SimpleQA as a non-math control to check that strong tail-reconstruction behavior does not appear uniformly across domains. While suffix reconstruction may be easier for templated math than for open-domain QA, the absence of a consistent advantage on SimpleQA suggests that the large Qwen deltas on legacy math are not explained by a domain-general completion skill alone.
@@
6 The Tax-Aware Minimum Standard for RLVR
Our results in section 4 and section 5 show that three pieces of evaluation controls materially change how RLVR gains appear: parity-controlled budgeting, calibration and attempt tracking, and contamination audits. Alongside synthesized prior work on judge robustness and safety evaluations, we package these into a tax-aware minimum standard for reporting RLVR results. The standard is intentionally narrow and measurement-focused. It does not prescribe a particular RL algorithm. Instead, it specifies the controls that must be in place before headline reasoning gains are treated as reliable.
(1) Budget parity and saturation curves.
First, RLVR models and their baselines must be evaluated under matched sampling budgets. In our gap analysis, several celebrated improvements shrank or disappeared once we fixed the verifier, prompt family, decoding parameters, and number of samples per item (Table 2). Reporting pass@k for the RLVR model and pass@1 for the base model, or silently changing temperature, top-$p$, or stopping rules, makes it impossible to tell how much of the gain comes from extra search rather than a better policy.
Under the minimum standard, any claim that RLVR improves a metric on a given benchmark must include: (i) the exact sampling budget for both base and RLVR models, (ii) a saturation curve that plots accuracy as a function of $k$ under a shared decoding setup, and (iii) at least three seeds with mean and confidence intervals or standard deviations (Mu et al., 2025; Hochlehnert et al., 2025). We recommend summarizing performance by area under the saturation curve in addition to a single pass@k point, since this is less sensitive to one particular choice of $k$ and reveals whether RLVR shifts the whole budget-performance frontier or only improves at very large $k$.
(2) Calibration, abstention, and judge robustness.
Second, evaluations must track calibration and abstention, not only accuracy. section 3 showed that RLVR often reduces refusals and increases stated confidence, which yields more attempted items but also more confident errors. Our factual QA control (Table 1) separates shared accuracy from newly attempted tail items and measures expected calibration error (ECE) on shared items. This reveals a distinct failure mode: headline scores rise because the model stops saying “I do not know” and starts answering everything, while accuracy on overlapping items barely moves.
Under the minimum standard, each reported accuracy figure must be accompanied by: (i) refusal or “not attempted” rates, (ii) shared accuracy on the intersection of items both models attempted, and (iii) a calibration metric such as ECE computed from the model’s confidence scores (Hochlehnert et al., 2025; Leng et al., 2024). For settings that rely on LLM-as-a-judge rather than programmatic verifiers, we also require at least one judge stress test: the same outputs scored under several prompt templates or instruction orderings, with the spread in scores reported (Zhao et al., 2025; Sun et al., 2025a; Yang et al., 2026). This does not remove judge fragility, but it makes visible how sensitive a claimed RLVR gain is to small changes in the judge pipeline. Together, these measurements ensure that gains are not driven primarily by attempt inflation, miscalibration, or a brittle judge configuration.
(3) Contamination audits and data hygiene.
Third, RLVR claims must be supported by explicit data provenance checks. section 5 used partial-prompt completion to show that some math benchmarks are heavily memorized by Qwen families: the model can reconstruct the hidden tail and final answer when given 80 percent of the problem on legacy sets, but not on fresh ones such as AIME-2025. Without such checks, apparent reasoning gains can simply reflect better recall of training data (Liu et al., 2024b; Liang et al., 2023; Wu and others, 2025).
Under the minimum standard, any benchmark used to support an RLVR claim must be accompanied by: (i) a contamination screen that combines fuzzy or lexical matching against pretraining and fine-tuning corpora (where available) and partial-prompt probes at several prefix lengths, and (ii) at least one clean held-out set that shows no evidence of tail reconstruction even at large prefixes. For benchmarks that do show contamination, we treat them as probes of memorization rather than as primary evidence of reasoning and clearly label them as such. We also require that dataset versions, prompt templates, and filtering rules be released or described precisely enough for others to re-run the audit.
Summary.
The tax-aware minimum standard asks for three things before treating RLVR gains as robust: matched budgets with saturation curves and variance disclosure, calibration and abstention metrics with at least one judge stress test when judges are used, and a contamination audit with at least one clean held-out benchmark. Taken together, these controls provide a simple, concrete bar that future RLVR work can meet without specifying a particular training recipe, and that can be reused for SFT, RLHF, and test-time compute evaluations.